OpenStudy (anonymous):

At a local flower shop, for every 13 bouquets ordered, 5 contain roses. If the flower shop sold 104 bouquets last month, what is a reasonable prediction of the number of bouquets that contained roses last month? A. 45 B. 39 C. 35 D. 40

OpenStudy (anonymous):

I'll medal and fan!!

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Could you help me with one more?

OpenStudy (anonymous):

Yes. but what I did was 104 * 5 / 13 that's how I got 40 if you want to make sure

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

A line has a y-intercept of 1 and a slope of 12/7. Find the equation of the line. A. y = x - 7/12 B. y = 7/12x - 1 C. y = x + 12/7 D. y = 12/7x + 1

OpenStudy (anonymous):

I think D am I right?

OpenStudy (anonymous):

I would also say its D

OpenStudy (anonymous):

Thank you. 2 more?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

What is the value of the 8th term in the sequence? -9 , -3 , 3 , 9 , 15 , ... A. 39 B. 33 C. 27 D. 45

OpenStudy (anonymous):

@Ciabella

OpenStudy (anonymous):

B. -9 , -3 , 3 , 9 , 15, 21 , 27 , 33

OpenStudy (anonymous):

Thank you !!!!

OpenStudy (anonymous):

welcome. don't you have one more ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Which of the following would be best displayed on a scatter plot? A. the percentage of restaurants that offer different brands of soda B. the amount of students signing up for a field trip C. the number of students wearing coats and the temperature outside D. the rate at which people walk based on the type of shoe they wear

OpenStudy (anonymous):

@Ciabella

OpenStudy (anonymous):

C:Two numerical variables - ideally suited for a scatter plot of coated students vs. temp.

OpenStudy (anonymous):

Thank you for all your help!!